OpenStudy (anonymous):

If h = 12 units and r = 5 units, then what is the volume of the cone shown above?

OpenStudy (k_lynn):

The volume formula of a cone is \[\frac{ 1 }{ 3 }\Pi r ^{2}h\] Now all you do is replace the variables into the equation. So \[\frac{ 1 }{ 3 }\times 3.14\times 5^{2}\times12\]

OpenStudy (k_lynn):

Can you multiply that for me @Icecube75789

OpenStudy (anonymous):

just a sec

OpenStudy (k_lynn):

okie doke.

OpenStudy (anonymous):

is it 100

OpenStudy (k_lynn):

You can find the decimal of the fraction 1 over 3 by dividing 1/3. When I punch this in my calculator, I get 0.333333..., then I multiply that by 3.14 and get 1.04666666..., the next part is multiplying that by \[5^{2}\] then I get 26.16666667. Now you just multiply that by 12, which is the height. Just plug the equation into your calculator.

OpenStudy (jordanloveangel):

hope this helped

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

yes